A wave composed of two sinusoidal waves of identical amplitude but different frequencies has the equation
s = 2A cos (45.0t) cos (67.0t)
where the coefficient of t is stated in rad/s.
(a) What is the lower of the two frequencies (in Hz) of the sinusoidal waves? (b) What is the higher of the two frequencies (in Hz) of the sinusoidal waves?
Answer #1

Given wave,

S=2Acos(45t)cos(67t)=A\left ( cos(45+67)t+cos(67-45)t \right )=Acos112t+Acos22t

a) So Lower frequency,

f_{L}=\frac{22}{2\pi}=3.5Hz

b) and higher one is,

f_{H}=\frac{112}{2\pi}=17.83Hz